How can I solve each system of equations? Example, x+y=-3 x-y=1 I think it has to do with substitution.
Found 2 solutions by Alan3354, LinnW:
Answer by Alan3354(69443)

x+y=-3
x-y=1
-------------- Add
2x = -2
etc
==============
You can use substitution.
---
x+y=-3
x-y=1 --> x = y+1
Sub for x in the 1st eqn
y+1 + y = -3
2y = -4
etc
-----
Elimination is the best for this example, imo.


Answer by LinnW(1048)

To use substitution we need to solve for x in one of the equations.
Let's take x-y=1
subtract y from each side
x = y + 1
The other equation is
x + y = -3
Substitute (y + 1) for x in this equation.
(y + 1) + y = -3
2y + 1 = -3
add -1 to each side
2y = -4
divide each side by 2
y = -2
Now substitute -2 for y in x = y + 1
x = -2 + 1
x = -1